To meet the grand challenge of climate change and reduce the rate of warming of the atmosphere we must begin to remove major sources of greenhouse gases from the atmosphere. One major greenhouse gas is methane. A major source of methane emissions are cows. There are approximately 1.4 billion cows (compare that to 7 billion people). Each cow emits approximately 120 kg/year of methane (CH4) Assume a cow’s stomach is 49 gallons in volume and the concentration in the stomach is 100 mg/L. Also assume the cow does not ingest or inhale any methane. Assume the stomach is well mixed, the cow does not ingest or inhale methane, there is no decay and no accumulation of methane in the stomach. (A) Write a full mass balance for methane from a cow and cross out the terms that you can eliminate based on the assumptions above and write the simplified mass balance (B) Calculate the average reaction rate of creation for methane in a cows stomach in units (per day). (B) How much methane (in kg) can be removed from the atmosphere per year if all cows are fed a diet or seaweed (or supplements) that reduce methane generation by 99%.
Answer #1

Part (b).

The average reaction rate of creation for methane in a cow's stomach = 120 kg/365 days = 0.329 kg/day or 329 g/day

Note: The overall rate of creation for methane from all cows' stomachs = 1.4*109*120 kg/year = 1.68*1011 kg/year

Part (c).

The methane that can be removed if all cows are fed a diet or seaweed (or supplements) that reduce methane generation by 99% = 1.4*109*120*99/100 = 1.6632*1011 kg/year